# Quantum Buffer Design Using Petri Nets

### v1

@@ 5 Concluding Remarks
This paper simplified the quantum Petri net (QPN) model given in [24] and used it to introduce quantum buffer designs to route quantum tokens between a set of input places and a set of output places. It presented a quantum S-R flip-flop that can be used to obtain quantum registers in order to hold quantum tokens at the input and output places and the places within the proposed quantum Petri net buffer designs. Ancillary $q$-tokens have been used in all five designs to fire transitions in order to route data $q$-tokens to their target (output) places. In particular, ancillary $q$-tokens were used as addresses either to select data $q$-tokens from the input places (MISO quantum buffer) or target output places to transfer $q$-tokens (SIMO quantum buffer) or both (MIMO quantum buffer). The design of the quantum S-R-flip-flop was validated using the IBM composer and running the composed quantum circuit on one of IBM’s quantum computers. The quantum buffer designs have been validated using Qiskit and Python. It is noted that, in the MIMO quantum buffer design, the placement of ancillary $q$-tokens into separate places forces a serialization of the transfer of data $q$-tokens from input places to output places. This is done for the clarity of presentation of the MIMO quantum buffer design. The serialization can easily be removed by integrating the ancillary $q$-tokens to the data $q$-tokens within the input places, in which case data $q$ tokens can be transferred from input places to output places in parallel according to any assignment between the input and output places. It should be added that a SISO quantum buffer can used to obtain FIFO, LIFO, and any other buffer with a single input place, single output place and a desired ordering of transfers of data $q$-tokens from the input place to the output place by replacing the ancillary qubits by $q$-tokens by the indicies of the data $q$-tokens in the input place. As for future work, it will be worthwhile to validate our quantum buffer designs on an actual quantum computer rather than by Qiskit simulations only. Another direction of research will be to extend the quantum Petri net buffer designs presented here to Petri net versions of multistage quantum packet switching networks[32, 33]. The exploration of these ideas will be deferred to another place.

### v2

@@ 6 Concluding Remarks
This paper simplified the quantum Petri net (QPN) model given in [25] and used it to introduce quantum buffer designs to route quantum tokens between a set of input places and a set of output places. It presented a quantum S-R flip-flop to design quantum registers to hold quantum tokens both at input and output places and the places within the proposed quantum buffer designs. Ancillary $q$-tokens have been used in all five quantum buffer designs to fire transitions in order to route data $q$-tokens to their target (output) places. In particular, ancillary $q$-tokens were used as addresses either to select data $q$-tokens from the input places (MISO quantum buffer) or target output places to transfer $q$-tokens (SIMO quantum buffer) or both (MIMO quantum buffer). Quantum buffers presented here are generalizations of classical buffers and as such they are expected to play a crucial role in quantum computers and networks much the same way classical buffers do in classical computers and networks. Thus, each type of quantum buffer plays a similar role in a quantum system as its counterpart does in the classical domain. For example, a quantum priority buffer may be used to prioritize the processing of some $q$-tokens over others because of the urgency of service they would need.
The quantum buffer designs have been validated using the IBM composer and quantum computers as well as Qiskit and Python to simulate the quantum buffer designs. It is noted that in the MIMO quantum buffer design, the placement of ancillary $q$-tokens into separate places forces a serialization of the transfer of data $q$-tokens from input places to output places. This is done for the clarity of presentation of the MIMO quantum buffer design. The serialization can easily be removed by integrating the ancillary $q$-tokens to the data $q$-tokens within the input places, in which case data $q$ tokens can be transferred from input places to output places in parallel according to any assignment between the input and output places. As for future work, it will be worthwhile to validate our quantum buffer designs on a quantum computer as our validations were carried out in Qiskit only. Another direction of investigation is to look into the use of our quantum buffer designs as a basis for constructing quantum packet switching networks [35, 36]. These and other further investigations will be deferred to another place.